OpenStudy (anonymous):

10x^2+21xy-10y^2/45x^2-23xy+2y^2/4x^2-20xy+25y^2/18x^2-47xy+5y^2 divide express the answer in lowest term

OpenStudy (tkhunny):

You will have to do better than that. I explain: a + b / c + d means \[a + \frac{b}{c} + d\] (a+b)/(c+d) means \[\frac{a+b}{c+d}\] I'm thinking you should be much more generous with parentheses or we can't tell what expression you intended.

OpenStudy (anonymous):

Thanks should i retype the problem im having trouble factoring

OpenStudy (tkhunny):

Yes, please. Add parentheses where you intend to group. Maybe: (10x^2+21xy-10y^2)/(45x^2-23xy+2y^2)/(4x^2-20xy+25y^2)/(18x^2-47xy+5y^2). I'm not sure. It's a pretty complicated expression. You'll have to be particularly clear.

OpenStudy (anonymous):

thats correct

OpenStudy (tkhunny):

Piece at a time: 10x^2+21xy-10y^2 = (5x-2y)*(2x-5y) Youdo the next one.